### v1

# MSHyper: Multi-Scale Hypergraph Transformer for Long-Range Time Series Forecasting

Abstract
[…]
1 Introduction
Many time series demonstrate complex and diverse temporal patterns of different scales Wen et al. (2021); Chen et al. (2022, 2023a); Zhang and Yan (2023). For example, due to periodic human activities, traffic occupation and electricity consumption show clear daily and weekly patterns. Considering interactions between these temporal patterns often leads to more accurate forecasting results than analyzing each pattern separately. For example, the morning rush hour on the first workday after a holiday tends to be more congested (interactions between daily and weekly patterns), while the evening rush hour on the last workday before a long holiday starts earlier. Therefore, how to model complex temporal patterns of different scales and their interactions is a fundamental problem in long-range time series forecasting.
To model temporal patterns of different scales and their interactions, traditional methods, e.g., seasonal ARIMA Box and Jenkins (1968) and Prophet Taylor and Letham (2018), use decomposition with heuristic priors to obtain temporal patterns of different scales, but cannot model complex non-linear dependencies of time series. Recently, deep neural networks have demonstrated superiority in capturing non-stationary and non-linear dependencies. Temporal convolutional networks (TCNs) Bai et al. (2018), recurrent neural networks (RNNs) Salinas et al. (2020), and Transformers Zhou et al. (2021); Wu et al. (2021) have been used for time series forecasting. To model temporal patterns of different scales, multi-scale Transformer-based methods Li et al. (2019) attempt to build sub-sequences of different scales from the original input sequence but ignore interactions between temporal patterns of different scales. To address this issue, recent multi-scale Transformer-based methods Liu et al. (2021); Cirstea et al. (2022) introduce special structures (e.g., pyramidal structures) between sub-sequences of different scales. These structures model temporal dependencies within a sub-sequence through intra-scale edges, and model interactions between temporal patterns of different scales through inter-scale edges.
However, these methods only use edges to model pairwise interactions and lack the ability to model high-order interactions (i.e., simultaneous interactions between multiple temporal patterns). In reality, temporal patterns of different scales co-exist and exhibit high-order interactions, e.g., the peak household electricity consumption during summer weekend afternoons (high-order interactions between daily, weekly, and monthly patterns), as well as the high but relatively stable household electricity consumption during winter weekends.
[…]
1) We propose a hypergraph and hyperedge graph construction (H-HGC) module that builds the hypergraph according to the temporal proximity rules, which can model intra-scale, inter-scale, and mixed-scale high-order interactions between temporal patterns. In addition, by treating hyperedges as nodes and building edges based on the sequential relationship and association relationship between nodes, H-HGC also builds the hyperedge graph to enhance hypergraph modeling.
2) We propose a tri-stage message passing (TMP) mechanism that has three message passing phases: Node-hyperedge, hyperedge-hyperedge, and hyperedge-node, which can aggregate pattern information and learn the interaction strength between temporal patterns of different scales.
3) We conduct extensive experiments on five real-world time series datasets, experimental results demonstrate that MSHyper achieves competitive performance, reducing prediction errors by an average of 8.73% and 7.15% over the best baseline in MSE and MAE, respectively.
2 Related Work
Methods for Time Series Forecasting. Time series forecasting methods can be roughly divided into statistical methods and deep neural network-based methods. Statistical methods (e.g., ARIMA Box and Jenkins (1968) and Prophet Taylor and Letham (2018)) follow arbitrary yet simple assumptions, and fail to model complicated temporal dependencies. Recently, deep neural networks show superiority in modeling complicated temporal dependencies. TAMS-RNNs Chen et al. (2021) obtains the periodic temporal dependencies through multi-scale recurrent neural networks (RNNs) with different update frequencies. LSTNet Lai et al. (2018) utilizes recurrent-skip connections in combination with convolutional neural networks (CNNs) to capture long- and short-term temporal dependencies. Transformer-based methods Wen et al. (2022); Chen et al. (2023b) take advantage of the attention mechanism and achieve great access in modeling long-range temporal dependencies. Reformer Kitaev et al. (2020) approximates the attention value through local-sensitive hashing (LSH), Informer Zhou et al. (2021) obtains the dominant query by calculating the KL-divergence to realize the approximate of self-attention, and Autoformer Wu et al. (2021) introduces an auto-correlation mechanism to operate at the level of sub-sequences. However, the above methods struggle to obtain the complex temporal patterns of long-range time series.
Multi-Scale Transformers. Multi-scale or hierarchical Transformers have been proposed in different fields, e.g., natural language processing Ainslie et al. (2020); Guo et al. (2019); Subramanian et al. (2020), computer version Zhang et al. (2021); Dosovitskiy et al. (2021); Wang et al. (2021), and time series forecasting Chen et al. (2022); Zhou et al. (2022); Shabani et al. (2022). Multi-scale ViT Zhang et al. (2021) realizes image recognition by combining multi-scale image feature embeddings and Transformer. Star-Transformer Guo et al. (2019) models intra-scale and inter-scale information interactions by introducing the global node embedding. Due to the limited expression capacity of a single global node, ETC Ainslie et al. (2020) carries out information interactions between the global and local nodes by introducing a set of global node embeddings and setting fixed-length windows. To further extend the ability to model interactions between temporal patterns of different scales, Pyraformer Liu et al. (2021) extends the two-layer structure into multi-scale embeddings, and models interactions between nodes of different scales through a pyramid graph. Crossformer Zhang and Yan (2023) combines a two-stage attention with a hierarchical encoder-decoder architecture to capture cross-time and cross-dimension interactions. However, existing methods only model pairwise interactions between nodes, ignoring high-order interactions between temporal patterns of different scales.
Hypergraph Neural Networks. Hypergraph neural networks (HGNNs) have been proven to be capable of modeling high-order interactions, which have been applied to various fields, e.g., visual object recognition Yan et al. (2020), trajectory prediction Xu et al. (2022), stock selection Sawhney et al. (2021), and citation network classification Bai et al. (2021). HGNN Feng et al. (2019) and HyperGCN Yadati et al. (2019) are the first works to apply graph convolution to hypergraphs, which demonstrate the superiority of hypergraphs over ordinary graph neural networks (GNNs) in modeling high-order interactions. Recent works Yang et al. (2020); Xia et al. (2022) show that HGNNs are promising to model pattern interactions from high-order relations. MBHT Yang et al. (2022) combines hypergraphs with a Transformer framework for the sequential recommendation, which leverages hypergraphs to capture high-order user-item interaction patterns. GroupNet Xu et al. (2022) uses multi-scale hypergraphs for trajectory prediction, which leverages topology inference and representation learning to capture the agent patterns with their high-order interactions.
Considering the ability of HGNNs in high-order interaction modeling, we propose a multi-scale hypergraph Transformer framework to model high-order interactions between temporal patterns of different scales. Specifically, a H-HGC module is introduced to build the hypergraph and hyperedge graph. In addition, a TMP mechanism is proposed to aggregate pattern information and learn the interaction strength between temporal patterns of different scales by three message passing phases.
3 Preliminaries
Hypergraph. A hypergraph is defined as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where =$\mathcal{E}=\{e_{1},e_{2},...,e_{M}\}$ is the hyperedge set and $\mathcal{V}=\{v_{1},v_{2},...,v_{N}\}$ is the node set. Each hyperedge $e_{m}\in\mathcal{E}$ contains a set of nodes $\{v_{1},v_{2},\ldots,v_{n}\}\subseteq\mathcal{V}$. The hypergraph $\mathcal{G}$ can be denoted as an incidence matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$, where $\mathbf{H}_{nm}=1$ if the $n\text{th}$ node belongs to the $m\text{th}$ hyperedge, else $\mathbf{H}_{nm}=0$. The degree of the $n\text{th}$ node is defined as follows:
$\mathbf{D}_{v_{n}}=\sum_{m=1}^{M}\mathbf{H}_{nm}.$ (1)
The degree of the $m\text{th}$ hyperedge is defined as follows:
$\mathbf{D}_{e_{m}}=\sum_{n=1}^{N}\mathbf{H}_{nm}.$ (2)
The results of node degrees and hyperedge degrees are stored in diagonal matrices $\textbf{D}_{\text{v}}\in\mathbb{R}^{N\times N}$ and $\mathbf{D}_{\mathrm{e}}\in\mathbb{R}^{M\times M}$, respectively.
Problem Statement. The task of long-range time series forecasting is to predict the future $H$ steps given the previous $T$ steps of observed values. Specifically, given the input sequence $\mathbb{X}_{1:T}^{\text{I}}=\left\{\bm{x}_{t}\mid\bm{x}_{t}\in\mathbb{R}^{D},t\in[1,T]\right\}$, where $\bm{x}_{t}$ represents the values at time step $t$, and $D$ is the feature dimension. The prediction task can be formulated as:
$\mathbb{X}_{T+1:T+H}^{\text{O}}=\mathcal{F}\left(\mathbb{X}_{1:T}^{\text{I}};\theta\right)\in\mathbb{R}^{H\times D},$ (3)
where $\mathbf{X}_{T+1:T+H}^{\text{O}}$ denotes the forecasting results, $\mathcal{F}$ denotes the mapping function, and $\theta$ denotes the learnable parameters of $\mathcal{F}$.
4 MSHyper
As mentioned above, the core of MSHyper is to build multi-scale hypergraph structures, which can explicitly model high-order interactions between temporal patterns of different scales. To accomplish this goal, we first map the input sequence into multi-scale embeddings through the multi-scale feature extraction (MFE) module and then leverage the H-HGC module to build the hypergraph and hyperedge graph. Finally, we employ the TMP mechanism to aggregate pattern information and learn the interaction strength between temporal patterns of different scales. Figure 1 illustrates the framework of MSHyper.
Figure 1: The framework of MSHyper, which consists of three parts: The multi-scale feature extraction (MFE) module, the hypergraph and hyperedge graph construction (H-HGC) module, and the tri-stage message passing (TMP) mechanism.
4.1 Multi-Scale Feature Extraction Module
To get the feature embeddings of different scales, we first map the input sequence into multi-scale embeddings. Concretely, suppose $\mathbf{X}^{s}=\{\bm{x}_{t}^{s}|\bm{x}_{t}^{s}\in\mathbb{R}^{D},t\in[1,h^{s}]\}$ denotes the sub-sequence at scale $s$, where $s=1,...,S$ denotes the scale index, and $S$ is the total number of scales. $h^{s}=\left\lfloor\frac{h^{s-1}}{l^{s-1}}\right\rfloor\text{s.t.}~{}{}s\geq 2$ is the horizon at scale $s$ and $l^{s-1}$ denotes the size of the aggregation window at scale $s-1$. $\mathbf{X}^{1}=\mathbf{X}^{\text{I}}_{1:T}$ is the raw input sequence and the aggregation process can be formulated as:
$\mathbf{X}^{s+1}=Aggregation(\mathbf{X}^{s};\theta^{s})\in\mathbb{R}^{h^{s+1}\times D},$ (4)
where $Aggregation$ is the aggregation function, e.g., 1D convolution or average pooling, and $\theta^{s}$ denotes the learnable parameters of the aggregation function at scale $s$.
4.2 Hypergraph and Hyperedge Graph Construction Module
4.2.1 Hypergraph Construction
Despite existing methods Liu et al. (2021); Wu et al. (2019) being capable of modeling interactions between temporal patterns of different scales through graph structures, we argue that these methods still face two limitations: (1) These methods fail to model high-order interactions between temporal patterns. (2) These methods limit interactions to the neighboring nodes and thus are incapable of capturing interactions between nodes that are far away but still show correlations. To address these problems, we build the hypergraph structures according to the temporal proximity rules. As shown in Figure 2, on the one hand, we build different types of hypergraphs to model intra-scale, inter-scale, and mixed-scale high-order interactions. On the other hand, we build the $k$-hop hypergraph under each type of hypergraph to aggregate information from different ranges of neighbors.
Figure 2: Hypergraph construction. (a), (b), and (c) represent the intra-scale hypergraph, inter-scale hypergraph, and mixed-scale hypergraph, respectively.
Intra-Scale Hypergraph. As shown in Figure 2(a), in order to capture high-order interactions between intra-scale temporal patterns, we construct the intra-scale hypergraph $\mathcal{G}_{\text{ita}}$, which contains two hypergraphs, i.e., the original intra-scale hypergraph $\mathcal{G}_{\text{o, ita}}$ and $k$-hop intra-scale hypergraph $\mathcal{G}_{\text{k, ita}}$. Mathematically, let $\mathcal{G}_{\text{o, ita}}=\{\mathcal{V},\mathcal{E}_{\text{o, ita}}\}$ be the original intra-scale hypergraph, where $\mathcal{E}_{\text{o, ita}}=\{\mathcal{E}_{\text{o, ita}}^{s}\}_{s\in\{1,...,S\}}$ contains original intra-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ $e_{i}^{s}\in\mathcal{E}_{\text{o, ita}}^{s}$ is defined as follows:
$e_{i}^{s}=\left\{v_{\epsilon}^{s},\forall v_{j}^{s}\in\mathcal{N}(v_{\epsilon}^{s})\right\}\text{s.t.}~{}{}0<j-\epsilon\leq H_{s},$ (5)
where $\epsilon=(i-1)H_{s}+1$ is the starting node index under the $i\text{th}$ hyperedge of scale $s$ based on original connections. $H_{s}$ is the number of nodes connected by each hyperedge of scale $s$ and $\mathcal{N}(v_{\epsilon}^{s})$ is the neighboring nodes connected to node $v_{\epsilon}^{s}$. Meanwhile, let $\mathcal{G}_{\text{k, ita}}=\{\mathcal{V},\mathcal{E}_{\text{k, ita}}\}$ be the $k$-hop intra-scale hypergraph, where $\mathcal{E}_{\text{k, ita}}=\{\mathcal{E}_{\text{k, ita}}^{s}\}_{s\in\{1,...,S\}}$ contains $k$-hop intra-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ based on $k$-hop connections $e^{s}_{i,k}\in\mathcal{E}^{s}_{\text{k, ita}}$ is defined as follows:
$e_{i,k}^{s}=\{v_{d}^{s},v_{d+k}^{s},\ldots,v_{d+(H_{s}-1)k}^{s}\},$ (6)
where $d$ is the starting node index under the $i\text{th}$ hyperedge of scale $s$ based on $k$-hop connections, which can be defined as follows:
$d=\left\lfloor\frac{i-1}{k}\right\rfloor\times H_{s}k+(i-1)\%k+1,$ (7)
where $k$ is the temporal distance between two neighboring nodes. The intra-scale hypergraph $\mathcal{G}_{\text{ita}}=\{\mathcal{V},\mathcal{E}_{\text{ita}}\}$ based on the original intra-scale hypergraph and the $k$-hop intra-scale hypergraph can be formulated as follows:
$\mathcal{G}_{\text{ita}}=concat(\mathcal{G}_{\text{o, ita}},\mathcal{G}_{\text{k, ita}}),$ (8)
where $concat$ denotes the concatenation operation. High-order interactions between temporal patterns not only exist between intra-scale temporal patterns, but also between inter-scale temporal patterns, e.g., interactions between hourly and daily patterns, and interactions between daily and weekly patterns. In addition, there are temporal pattern interactions across all scales, e.g., interactions between hourly, daily, and weekly patterns. Therefore, we design the inter-scale hypergraph and mixed-scale hypergraph.
Inter-Scale Hypergraph. As shown in Figure 2(b), the inter-scale hypergraph $\mathcal{G}_{\text{ite}}=\{\mathcal{V},\mathcal{E}_{\text{ite}}\}$ is obtained by concatenating the original inter-scale hypergraph and $k$-hop inter-scale hypergraph. Mathematically, let $\mathcal{G}_{\text{o, ite}}=\{\mathcal{V},\mathcal{E}_{\text{o, ite}}\}$ be the original inter-scale hypergraph, where $\mathcal{E}_{\text{o, ite}}=\{\mathcal{E}_{\text{o, ite}}^{s}\}_{s\in\{1,...,S\}}$ contains original inter-scale hyperedge sets of different scales. The $i\text{th}$ hyperedge of scale $s$ $e^{s}_{i}\in\mathcal{E}^{s}_{\text{o, ite}}$ is defined as follows:
$e_{i}^{s}=\left\{v_{\lceil\epsilon/{l^{s}}\rceil}^{s+1},v_{\epsilon}^{s},v_{\epsilon+1}^{s},\ldots,v_{\epsilon+H_{s}}^{s}\right\},$ (9)
where $l^{s}$ denotes the size of the aggregation window at scale $s$. Then, let $\mathcal{G}_{\text{k, ite}}=\{\mathcal{V},\mathcal{E}_{\text{k, ite}}\}$ be the $k$-hop inter-scale hypergraph, where $\mathcal{E}_{\text{k, ite}}=\{\mathcal{E}_{\text{k, ite}}^{s}\}_{s\in\{1,...,S\}}$ contains $k$-hop inter-scale hyperedge sets of different scales and $e^{s}_{i,k}\in\mathcal{E}^{s}_{\text{k, ite}}$ is defined as follows:
$e^{s}_{i,k}=\{v_{\lceil d/l^{s}\rceil}^{s+1},v_{d}^{s},v_{d+k}^{s},\dots,v_{d+(H_{s}-1)k}^{s}\}.$ (10)
Mixed-Scale Hypergraph. As shown in Figure 2(c), the mixed-scale hypergraph $\mathcal{G}_{\text{mix}}=\{\mathcal{V},\mathcal{E}_{\text{mix}}\}$ is obtained by concatenating the original mixed-scale hypergraph and $k$-hop mixed-scale hypergraph. Mathematically, let $\mathcal{G}_{\text{o, mix}}=\{\mathcal{V},\mathcal{E}_{\text{o, mix}}\}$ be the original mixed-scale hypergraph, where $\mathcal{E}_{\text{o, mix}}$ denotes the original mixed-scale hyperedge set. The $i\text{th}$ hyperedge $e_{i}\in\mathcal{E}_{\text{o, mix}}$ is defined as follows:
$e_{i}=\left\{v_{\delta_{S}}^{S},v_{\delta_{S-1}}^{S-1},\dots,v_{\delta_{1}}^{1},v_{\delta_{1}+1}^{1},\dots,v_{\delta_{1}+H-1}^{1}\right\},$ (11)
where $\delta_{s}=\lceil\epsilon/(1\times\prod_{\alpha=2}^{s}l^{\alpha})\rceil$ is the starting node index of scale $s$ under the $i$th hyperedge based on original connections. Then, let $\mathcal{G}_{\text{k, mix}}=\{\mathcal{V},\mathcal{E}_{\text{k, mix}}\}$ be the $k$-hop mixed-scale hypergraph, where $\mathcal{E}_{\text{k, mix}}$ denotes the $k$-hop mixed-scale hyperedge set. $e_{i,k}\in\mathcal{E}_{\text{k, mix}}$ is defined as follows:
[…]
where $\delta_{s,k}=\lceil{d/(1\times\prod_{\alpha=2}^{s}l^{\alpha})}\rceil$ is the starting node index of scale $s$ under the $i$th hyperedge based on $k$-hop connections.
The final hypergraph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ is obtained by considering the abovementioned three hypergraphs, which can be formulated as follows:
$\mathcal{G}=concat(\mathcal{G}_{\text{ita}},\mathcal{G}_{\text{ite}},\mathcal{G}_{\text{mix}}).$ (13)
4.2.2 Hyperedge Graph Construction
After building the hypergraph, to enhance hypergraph modeling, we build the hyperedge graph to model interactions between hyperedges. Specifically, we build the hyperedge graph by representing hyperedges in $\mathcal{G}$ as nodes and considering the sequential relationship and association relationship. Let $G=(V,E,\mathbf{A})$ be the hyperedge graph, where $V=\{v_{e_{i}}|v_{e_{i}}\in\mathcal{E}\}$ is the node set of $G$ and $E=\{(v_{e_{i}},v_{e_{j}})|v_{e_{i}},v_{e_{j}}\in\mathcal{E}\}$ is the edge set of $G$. $\mathbf{A}$ is the weighted adjacency matrix defined based on the sequential relationship and association relationship between hyperedges.
Sequential Relationship. Within the intra-scale hypergraph, two hyperedges that connect nodes with temporal order exhibit the sequential relationship. For example, as shown in Figure 3(a), hyperedge $e_{1}$ connects the previous three nodes, and hyperedge $e_{2}$ connects the next three nodes. Since long-range time series is a collection of data points arranged in chronological order, intuitively, changes (e.g., increase, decrease, or fluctuation) in the values of nodes connected by $e_{1}$ may influence the values of nodes connected by $e_{2}$. Formally, let $G_{\text{sr}}=(V_{\text{sr}},E_{\text{sr}},\mathbf{A}_{\text{sr}})$ represent the hyperedge graph constructed based on the sequential relationship, where $V_{\text{sr}}=\{v_{{e}_{\text{sr}}}|v_{{e}_{\text{sr}}}\in\mathcal{E}_{\text{ita}}\}\in\mathbb{R}^{{D}_{\text{sr}}\times{D}}$ is the node set and $D_{\text{sr}}$ is the number of intra-scale hyperedges. $E_{\text{sr}}=\{(v_{{e}_{i,\text{sr}}},v_{{e}_{j,\text{sr}}})|v_{{e}_{i,\text{sr}}},v_{{e}_{j,\text{sr}}}\in\mathcal{E}_{\text{ita}},0\leq i-j\leq 1\}$ denotes the edge set constructed based on the sequential relationship. As shown in Figure 3(c), the adjacency matrix constructed based on the sequential relationship $\mathbf{A}_{\text{sr}}\in\mathbb{R}^{D_{\text{sr}}\times D_{\text{sr}}}$ is defined as follows:
[…]
Figure 3: Hyperedge graph construction. (a) and (b) are the sequential relationship and association relationship, respectively. (c) and (d) are the constructed adjacency matrix based on the sequential relationship and association relationship, respectively.
Association Relationship. Within the inter-scale hypergraph and mixed-scale hypergraph, we define hyperedges with common nodes have an association relationship. As shown in Figure 3(b), $e_{5}$ and $e_{7}$ have an association relationship, as they share a common node $n_{13}$. Formally, let $G_{\text{ar}}=(V_{\text{ar}},E_{\text{ar}},\mathbf{A}_{\text{ar}})$ represent the hyperedge graph constructed based on the association relationship, where $V_{\text{ar}}=\{v_{e_{\text{ar}}}|v_{e_{\text{ar}}}\in\{\mathcal{E}_{\text{ite}},\mathcal{E}_{\text{mix}}\}\}\in\mathbb{R}^{D_{\text{ar}}\times D}$ is the node set of $G_{\text{ar}}$ and $D_{\text{ar}}$ is the number sum of inter-scale hyperedges and mixed-scale hyperedges. $E_{\text{ar}}=\{(v_{{e}_{i,\text{ar}}},v_{{e}_{j,\text{ar}}})|v_{{e}_{i,\text{ar}}},v_{{e}_{j,\text{ar}}}\in\{\mathcal{E}_{\text{ite}},\mathcal{E}_{\text{mix}}\},|v_{{e}_{i,\text{ar}}}\cap v_{{e}_{j,\text{ar}}}|\geq 1\}$ denotes the edge set of $G_{\text{ar}}$. As shown in Figure 3(d), the adjacency matrix based on the association relationship $\mathbf{A}_{\text{ar}}\in\mathbb{R}^{D_{\text{ar}}\times D_{\text{ar}}}$ is defined as follows:
[…]
The adjacency matrix $\mathbf{A}$ based on the above two relationships can be defined as follows:
$\mathbf{A}=\begin{bmatrix}\mathbf{A}_{\text{sr}}&\mathbf{\Gamma}_{1}\\ \mathbf{\Gamma}_{2}&\mathbf{A}_{\text{ar}}\end{bmatrix}\in\mathbb{R}^{D_{\varphi}\times D_{\varphi}},$ (16)
where $\bm{\Gamma}_{1}\in\mathbb{R}^{D_{\text{sr}}\times D_{\text{ar}}}$ and $\bm{\Gamma}_{2}\in\mathbb{R}^{D_{\text{ar}}\times D_{\text{sr}}}$ are matrices consisting entirely of zeros, and $D_{\varphi}$ is the sum of ${D}_{\mathrm{\text{ar}}}$ and ${D}_{\mathrm{\text{sr}}}$.
4.3 Tri-Stage Message Passing Mechanism
After building the hypergraph and hyperedge graph, to aggregate pattern information and learn the interaction strength between temporal patterns of different scales, we propose a TMP mechanism, which contains the node-hyperedge, hyperedge-hyperedge, and hyperedge-node phases.
Node-Hyperedge Phase. Given the sequences based on the MFE module $\mathbf{X}=\{\mathbf{X}^{1},\mathbf{X}^{2},\ldots,\mathbf{X}^{S}\}\in\mathbb{R}^{N\times D}$, where $N$ denotes the number sum of input time steps and aggregated feature values of different scales. We first get the initialized node embeddings $\bm{\mathcal{V}}=f(\mathbf{X})\in\mathbb{R}^{N\times D}$, where $f$ can be implemented by the multi-layer perceptron (MLP). As shown in Figure 1(c), we get the initialized hyperedge embeddings by the aggregation operation based on the hypergraph $\mathcal{G}$. Specifically, for the $i\text{th}$ hyperedge $e_{i}\in\mathcal{E}$, its initialized embedding is obtained as follows:
$\bm{v}_{e_{i}}=\sum\limits_{v_{j}\in\mathcal{N}(e_{i})}\bm{v}_{j}\in\mathbb{R}^{D},$ (17)
where $\mathcal{N}(e_{i})$ denotes neighboring nodes connected by $e_{i}$.
Hyperedge-Hyperedge Phase. After getting initialized hyperedge embeddings, we proceed to update their embeddings through the constructed hyperedge graph. Specifically, for the given initialized hyperedge embeddings $\mathbf{V}=\{\bm{v}_{e_{i}}|\bm{v}_{e_{i}}\in\mathcal{E}\}\in\mathbb{R}^{M\times D}$, we transform it into query $\widetilde{\mathbf{Q}}=\mathbf{VW}^{\text{q}}$, key $\widetilde{\mathbf{K}}=\mathbf{V}\mathbf{W}^{\text{k}}$, and value $\widetilde{\mathbf{V}}=\mathbf{V}\mathbf{W}^{\text{v}}$, where $\mathbf{W}^{\text{q}}$, $\mathbf{W}^{\text{k}}$, and $\mathbf{W}^{\text{v}}$ are learnable weight matrices. For the $i\text{th}$ row $\bm{\widetilde{q}}_{i}\in\widetilde{\mathbf{Q}}$, the updated hyperedge embedding $\widetilde{\bm{v}}_{e_{i}}\in\widetilde{\mathbf{V}}_{\text{e}}$ can be calculated as follows:
$\widetilde{\bm{v}}_{e_{i}}=\sum_{j=1}^{M}\frac{exp(e_{ij})}{\Sigma_{\ell=1}^{M}exp(e_{i\ell})}\widetilde{\bm{v}}_{j}$ (18)
$e_{ij}=\frac{\widetilde{\bm{q}}_{i}\widetilde{\bm{k}}_{j}^{T}}{\sqrt{D}}-(1-\mathbf{A}_{ij})\text{C},$
where $\bm{\widetilde{k}}_{j}^{T}$ denotes the transpose of the $j\text{th}$ row in $\widetilde{\mathbf{K}}$ and $\bm{\widetilde{v}}_{j}$ denotes the $j\text{th}$ row in $\widetilde{\mathbf{V}}$. $\mathbf{A}_{ij}\in\mathbf{A}$ is a binary value and C is a large constant.
Hyperedge-Node Phase. After obtaining the updated hyperedge embeddings, we update the node embeddings by considering all the related hyperedges. Considering the constructed hypergraph $\mathcal{G}$, we use the hypergraph convolution to update the node embeddings. Specifically, the symmetric normalized hypergraph Laplacian convolution can be formulated as follows:
$\bm{\widetilde{\mathcal{V}}}=\sigma(\mathbf{D}_{v}^{-1/2}\mathbf{HD}_{e}^{-1}\mathbf{H}^{T}\mathbf{D}_{v}^{-1/2}\bm{{\mathcal{V}}}\mathbf{{P}})\in\mathbb{R}^{{N}\times{D}_{c}},$ (19)
where $\bm{\widetilde{\mathcal{V}}}$ is the output of the hypergraph convolution and ${D}_{c}$ is the output dimension. $\mathbf{P}\in\mathbb{R}^{D\times D_{c}}$ denotes the learnable parameters and $\sigma$ is the activation function, e.g., LeakyReLU and ELU. To capture the interaction strength of each node $v_{i}\in\bm{\mathcal{V}}$ and its related hyperedges, we dynamically update $\mathbf{H}$ using the node embedding and updated hyperedge embeddings, which can be formulated as follows:
$\mathbf{H}_{ij}^{\text{att}}=\frac{exp(\sigma(f_{t}[\bm{v}_{i},\widetilde{\bm{v}}_{e_{j}}]))}{\sum_{k\in\mathcal{N}_{i}}exp(\sigma(f_{t}[\bm{v}_{i},\widetilde{\bm{v}}_{e_{k}}])))},$ (20)
where [.,.][.,.][ . , . ] denotes the concatenation operation of the node and its related hyperedges. $f_{t}$ is a trainable MLP, and $\mathcal{N}_{i}$ is the neighboring hyperedges connected to $v_{i}$, which can be accessed using the constructed hypergraph $\mathcal{G}$. Then, we use Equation 19 to aggregate pattern information of different scales by replacing $\mathbf{H}$ with $\mathbf{H}^{\text{att}}$. The multi-head attention mechanism is also used to stabilize the training process:
$\bm{\widetilde{\mathcal{V}}}=\bigoplus_{\jmath=1}^{\mathcal{J}}(\sigma(\mathbf{D}_{v}^{-1/2}\mathbf{H}_{\jmath}^{\mathrm{att}}\mathbf{D}_{e}^{-1}\mathbf{H}_{\jmath}^{\mathrm{att}^{T}}\mathbf{D}_{v}^{-1/2}\bm{\mathcal{V}}\mathbf{P}_{\jmath})),$ (21)
where $\bigoplus$ is the aggregation function used for combing the outputs of multi-head, e.g., concatenation or average pooling. $\mathbf{H}^{\text{att}}_{\jmath}$ and $\mathbf{P}_{\jmath}$ are the enriched incidence matrix and the learnable weight matrix of the $j\text{th}$ head, respectively. $\mathcal{J}$ is the number of heads.
4.4 Objective Function
After obtaining the updated node embeddings encoded by the multi-scale hypergraph, we concatenate the last node embeddings of each sub-sequence from different scales and then put them into a linear layer for prediction. We choose MSE as our objective function, which is defined as follows:
$\mathcal{L}=\frac{1}{H}\left\|\widehat{\mathbf{X}}_{T+1:T+H}^{\text{O}}-\mathbf{X}_{T+1:T+H}^{\text{O}}\right\|_{2}^{2},$ (22)
where $\widehat{\mathbf{X}}_{T+1:T+H}^{\text{O}}$ and $\mathbf{X}_{T+1:T+H}^{\text{O}}$ are ground truth and forecasting results, respectively.
5 Experiments
5.1 Experimental Setup
Table 1: Multivariate long-range time series forecasting results on five real-world datasets. The input length is set as $I=168$, and the prediction length $O$ is set as 96, 168, 336, and 720 (For ETTm1, the prediction length is set as 96, 288, 672, and 720). IMP shows the improvement of MSHyper over the best baseline. The best results are bolded and the second best results are underlined.
Methods
Informer
(AAAI 2021)
Autoformer
(NeurIPS 2021)
Pyraformer
(ICLR 2021)
FEDformer
(ICML 2022)
Quatformer
(KDD 2022)
Crossformer
(ICLR 2023)
MSHyper
IMP
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather 96 0.300 0.384 0.266 0.336 0.503* 0.518* 0.217 0.296 0.211 0.279 0.444* 0.473* 0.164 0.217 22.27% 22.22%
168 0.608 0.567 0.574 0.548 0.519 0.521 0.564 0.541 0.251* 0.295* 0.473 0.494 0.197 0.246 21.51% 16.61%
336 0.702 0.620 0.600 0.571 0.539 0.543 0.533 0.536 0.310 0.344 0.495 0.515 0.262 0.298 15.48% 13.37%
720 0.831 0.731 0.587 0.570 0.547 0.553 0.562 0.557 0.381 0.374 0.526 0.542 0.309 0.325 18.90% 13.10%
Electricity 96 0.274 0.368 0.255 0.339 0.272* 0.371* 0.183 0.297 0.197 0.308 0.185* 0.278* 0.183 0.271 0% 2.52%
168 0.368 0.424 0.299 0.387 0.452 0.455 0.263 0.361 0.205* 0.315* 0.231 0.309 0.193 0.298 5.85% 3.56%
336 0.381 0.431 0.375 0.428 0.463 0.456 0.305 0.386 0.220 0.329 0.323 0.369 0.205 0.310 6.82% 5.78%
720 0.406 0.443 0.377 0.434 0.480 0.461 0.372 0.434 0.245 0.350 0.404 0.423 0.230 0.329 6.12% 6.00%
ETTh1 96 0.865 0.713 0.449 0.459 0.701* 0.635* 0.376 0.419 0.422* 0.447* 0.396* 0.412* 0.375 0.398 0.27% 3.40%
168 0.931 0.752 0.493 0.479 0.781 0.675 0.412 0.449 0.453 0.489 0.410 0.441 0.405 0.427 1.22% 3.17%
336 1.128 0.873 0.509 0.492 0.912 0.747 0.456 0.474 0.491 0.495 0.440 0.461 0.428 0.439 2.73% 4.77%
720 1.215 0.896 0.539 0.537 0.993 0.792 0.521 0.515 0.523 0.525 0.519 0.524 0.503 0.507 3.08% 1.55%
ETTm1 96 0.678 0.614 0.502 0.476 0.520 0.504 0.366 0.412 0.375 0.398 0.320 0.373 0.315 0.365 1.56% 2.14%
288 1.056 0.786 0.604 0.522 0.729 0.657 0.398 0.433 0.408 0.424 0.404 0.427 0.394 0.419 1.01% 1.17%
672 1.192 0.926 0.607 0.530 0.980 0.678 0.455 0.464 0.472* 0.468* 0.569 0.528 0.451 0.462 0.88% 0.43%
720 1.166 0.823 0.671 0.561 0.912* 0.727* 0.543 0.490 0.499 0.502 0.751* 0.677* 0.493 0.471 1.20% 3.88%
Traffic 96 0.719 0.391 0.613 0.388 0.628* 0.354* 0.587 0.366 0.618 0.384 0.450* 0.477* 0.408 0.269 9.33% 24.01%
168 0.660 0.391 0.649 0.407 0.635 0.347 0.607 0.385 0.621* 0.399* 0.513 0.289 0.418 0.271 18.52% 6.23%
336 0.747 0.405 0.624 0.388 0.641 0.347 0.624 0.389 0.622 0.384 0.530 0.300 0.431 0.287 18.68% 4.33%
720 0.792 0.430 0.674 0.417 0.670 0.364 0.623 0.378 0.629 0.383 0.573 0.313 0.463 0.298 19.20% 4.79%
  • * indicates that some methods do not have uniform prediction lengths with other methods. To ensure a fair comparison, we utilize their official code and adjust prediction lengths.
Datasets. We conduct experiments on five commonly used long-range time series forecasting datasets, including ETT (ETTh and ETTm), Electricity, Weather, and Traffic datasets. The descriptions about the five datasets are as follows:
ETT Zhou et al. (2021): This dataset contains the oil temperature and load data collected by electricity transformers, including ETTh (ETTh1 and ETTh2) and ETTm (ETTm1 and ETTm2), which are sampled hourly and every 15 minutes, respectively.
Electricity  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014: This dataset contains the electricity consumption of 321 clients from the UCI Machine Learning Repository, which is sampled hourly.
Traffic222 http://pems.dot.ca.gov: This dataset contains the road occupancy rates of 862 sensors in San Francisco Bay Area freeways, which is sampled hourly.
Weather333 https://www.bgc-jena.mpg.de/wetter/: This dataset contains 21 meteorological measurements data from the Weather Station of the Max Planck Biogeochemistry, which is sampled every 10 minutes.
Following existing works Zhou et al. (2021); Zhang and Yan (2023); Liu et al. (2021), we split each dataset into training, validation, and testing set based on chronological order. The ratio is 6:2:2 for the ETT dataset and 7:2:1 for the others.
Experimental Settings. MSHyper is trained/tested on a single NVIDIA Geforce RTX 3090 GPU. The aggregation windows are set to 4 for ETTh and ETTm datasets and 3 for other datasets. $S$ and $H$ in MSHyper are set to 4 in all experiments. Z-Score normalization is used to normalize all datasets. MSE and MAE are used as evaluation metrics. Lower MSE and MAE results mean better performance. Adam is set as the optimizer with the initial learning rate of $\text{10}^{\text{-4}}$, and the batch size is set to 32.
Baselines. The baselines include six latest Transformer-based models: Informer Zhou et al. (2021), Autoformer Wu et al. (2021), Pyraformer Liu et al. (2021), FEDformer Zhou et al. (2022), Quatformer Chen et al. (2022), and Crossformer Zhang and Yan (2023). The source code of MSHyper is released on Anonymous GitHub [^4] https://anonymous.4open.science/r/MSHyper-A3FF.
5.2 Main Results
Multivariate Results. Table 1 shows the multivariate experimental results of MSHyper compared with baselines on five datasets. We can see that MSHyper achieves state-of-the-art (SOTA) results on all five datasets at all prediction lengths. From Table 1 we can discern the following tendencies: 1) Informer and Autoformer exhibit relatively poor predictive performance. This may come from that time series forecasting requires modeling temporal pattern interactions, while vanilla attention or simplistic decomposition techniques are insufficient in capturing such interactions. 2) By considering the temporal pattern interactions of different scales, Pyraformer, Crossformer, FEDformer, and Quatformer all achieve excellent performance. 3) Benefiting from its ability to model high-order temporal pattern interactions, MSHyper achieves the best performance in long-range time series forecasting. Specifically, MSHyper reduces prediction errors by an average of 8.73% and 7.15% over the best baseline in MSE and MAE, respectively.
Univariate Results. Table 2 summaries the results of univariate long-range time series forecasting on ETTh (ETTh1 and ETTh2) and ETTm (ETTm1 and ETTm2) datasets. We can see from Table 2 that MSHyper achieves SOTA results on all datasets at all prediction lengths. Specifically, MSHyper gives an average error reduction of 9.57% and 7.09% compared to the best baseline in MSE and MAE, respectively. The experimental results verify the effectiveness of MSHyper in univariate long-range time series forecasting.
Table 2: Univariate long-range time series forecasting results on ETT dataset. The input length is set as $I$=168, and the prediction length $O$ is set as 96, 192, 336, and 720. IMP shows the improvement of MSHyper over the best baseline. The best results are bolded and the second best results are underlined.
Methods
Informer
(AAAI 2021)
Autoformer
Pyraformer*
(ICLR 2021)
FEDformer
(ICML 2022)
Quatformer*
(KDD 2022)
Crossformer*
(ICLR 2023)
MSHyper
IMP
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.193 0.377 0.071 0.206 0.099 0.277 0.079 0.215 0.069 0.222 0.076 0.216 0.060 0.187 13.04% 9.22%
192 0.217 0.395 0.114 0.262 0.174 0.346 0.104 0.245 0.138 0.222 0.085 0.225 0.078 0.215 8.24% 3.15%
336 0.202 0.381 0.107 0.258 0.198 0.370 0.119 0.270 0.123 0.275 0.106 0.257 0.087 0.226 17.92% 12.06%
720 0.183 0.355 0.126 0.283 0.209 0.384 0.142 0.299 0.146 0.293 0.128 0.287 0.108 0.230 14.29% 18.73%
ETTh2 96 0.213 0.373 0.153 0.306 0.152 0.303 0.128 0.271 0.135 0.278 0.125 0.273 0.117 0.266 6.40% 1.85%
192 0.227 0.387 0.204 0.351 0.197 0.348 0.185 0.330 0.189 0.342 0.187 0.334 0.172 0.325 7.03% 1.52%
336 0.242 0.401 0.246 0.389 0.238 0.385 0.231 0.378 0.234 0.375 0.227 0.377 0.222 0.373 2.20% 0.53%
720 0.291 0.439 0.268 0.409 0.274 0.435 0.278 0.420 0.263 0.424 0.266 0.410 0.242 0.401 7.98% 1.96%
ETTm1 96 0.109 0.277 0.056 0.183 0.127 0.281 0.033 0.140 0.035 0.147 0.035 0.145 0.032 0.136 3.03% 2.86%
192 0.151 0.310 0.081 0.216 0.205 0.343 0.058 0.186 0.174 0.193 0.055 0.180 0.049 0.168 10.91% 6.67%
336 0.427 0.591 0.076 0.218 0.302 0.457 0.084 0.231 0.105 0.237 0.072 0.209 0.066 0.195 8.33% 6.70%
720 0.438 0.586 0.110 0.267 0.387 0.485 0.102 0.250 0.112 0.264 0.097 0.248 0.087 0.224 10.31% 9.68%
ETTm2 96 0.088 0.225 0.065 0.189 0.074 0.208 0.067 0.198 0.075 0.217 0.058 0.183 0.047 0.165 18.97% 9.84%
192 0.132 0.283 0.118 0.256 0.116 0.252 0.102 0.245 0.117 0.263 0.105 0.237 0.094 0.212 7.84% 10.55%
336 0.180 0.336 0.154 0.305 0.143 0.295 0.130 0.279 0.145 0.286 0.133 0.280 0.123 0.242 5.38% 13.26%
720 0.300 0.435 0.182 0.335 0.197 0.338 0.178 0.325 0.183 0.368 0.181 0.324 0.158 0.308 11.24% 4.94%
  • * indicates that some methods do not yield results for univariate long-range time series forecasting. To ensure a fair comparison, we utilize their official code and fine-tune key hyperparameters via grid search. Other results are from FEDformer.
5.3 Ablation Studies
We conduct ablation studies to verify the impact of different components on long-range time series forecasting. All ablation studies are conducted on ETTh1 dataset.
Multi-Scale Hypergraph. To investigate the effectiveness of the multi-scale hypergraph, we conduct ablation studies by carefully designing the following three variants: 1) Replacing the H-HGC model with the fully-connected graph (MSHyper-FCG). 2) Replacing the H-HGC model with the pyramid graph used in Liu et al. (2021) (MSHyper-PG). 3) Connecting the input sequence with intra-scale hyperedges, and thus the multi-scale hypergraph turns to the single-scale hypergraph (MSHyper-SSH). The experimental results are shown in Table 3, from which we can observe that MSHyper performs the best in all cases, which indicates the importance of multi-scale hypergraph in modeling high-order interactions between temporal patterns of different scales.
Table 3: Results of different multi-scale hypergraph construction methods.
Methods Prediction length
168 336 720
MSHyper-FCG
MSE
MAE
1.035
0.863
1.079
0.785
1.247
0.896
MSHyper-PG
MSE
MAE
0.712
0.649
0.882
0.725
0.892
0.747
MSHyper-SSH
MSE
MAE
0.601
0.568
0.643
0.587
0.674
0.629
MSHyper
MSE
MAE
0.405
0.427
0.428
0.439
0.503
0.507
Hyperedge Graph. To investigate the effectiveness of the hyperedge graph, we conduct ablation studies by carefully designing the following three variants: 1) Removing the association relationship of the hyperedge graph (MSHyper-w/o AR). 2) Removing the sequential relationship of the hyperedge graph (MSHyper-w/o SR). 3) Removing the hyperedge graph, i.e., without the hyperedge-hyperedge phase (MSHyper-w/o H-H). The experimental results are shown in Table 4. We can observe that MSHyper performs better than MSHyper-w/o AR and MSHyper-w/o SR, showing the effectiveness of the association relationship and sequential relationship, respectively. In addition, removing the hyperedge-hyperedge phase gets the worst forecasting results, which demonstrates the superiority of the hyperedge graph in enhancing hypergraph modeling.
Table 4: Results of different hyperedge graph construction methods.
Methods Prediction length
168 336 720
MSHyper-w/o AR
MSE
MAE
0.413
0.429
0.438
0.445
0.514
0.505
MSHyper-w/o SR
MSE
MAE
0.408
0.419
0.433
0.452
0.507
0.516
MSHyper-w/o H-H
MSE
MAE
0.430
0.474
0.531
0.506
0.612
0.598
MSHyper
MSE
MAE
0.405
0.427
0.428
0.439
0.503
0.507
5.4 Parameter Studies
We also perform parameter studies to measure the impact of the aggregation function, $k$-hop, and input length on ETTh1 dataset. The experimental results are shown in Table 5, Figure 5, and Figure 5, from which we can observe that: 1) The results based on the convolution operation outperform those based on other operations, possibly because convolution kernels can extract more complex features. 2) The best performance can be obtained when the $k$-hop is 4. The reason may be that a small $k$-hop limits interactions to the neighboring nodes, and a large $k$-hop would introduce noises. 3) MSHyper achieves the optimal performance when the input length is 168. The reason may be that the sequence with a shorter input length contains insufficient pattern information, while the sequence with a longer input length introduces more noise interference for high-order pattern interactions.
Table 5: Impact of aggregation function.
Methods Prediction length
168 336 720
Max-pooling
MSE
MAE
0.508
0.468
0.437
0.457
0.513
0.524
Average-pooling
MSE
MAE
0.532
0.506
0.431
0.455
0.512
0.523
Conv
MSE
MAE
0.405
0.427
0.428
0.439
0.503
0.507
Figure 4: Impact of $k$-hop.
Refer to caption k-hop
MSE
Refer to caption Input length
MSE
Figure 4: Impact of $k$-hop.
Figure 5: Impact of input length.
6 Conclusions
In this paper, we propose MSHyper for long-range time series forecasting. Specifically, the H-HGC module is introduced to provide foundations for modeling high-order interactions between temporal patterns. The TMP mechanism is employed to aggregate high-order pattern information and learn the interaction strength between temporal patterns of different scales. Extensive experiments on five real-world datasets show the superiority of MSHyper.